OpenStudy (anonymous):

a company had a total debt of $320,000 in 1980. between 1980 and 1987 it was able to reduce its debt 15% each year. approximate the companys debt in 1987 to the nearest 1000

OpenStudy (gw2011):

The answer to the nearest 1000 is $103,000 which is for the end of 1987 . Here is how you can calculate it: (D)(0.85)^7 where D = the total debt (0.85)^7 = the amount of debt remaining which is taken to the power of 7 since the debt is reduced each year beginning in 1981 and going to the end of 1987 which is 7 years. Therefore, the formula can now be used as follows: (D)(0.85)^7 = ($320,000)(0.85)^7 = $102,584.67 and to the nearest 1000 the answer is $103,000. I've assumed that the total debt at the end of 1980 is $320,000 and that the debt is to be calculated for the end of 1987.